

********************************************************************
UARS WEEKLY STATUS REPORT - 06 July 2001
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM did not collect data for June 27 through July 1. SUSIM was
repowered on July 1, but could not collect data because the instrument
had become too cold and required an extra day to warm up.
Afterward, SUSIM has completed its daily solar scans during periods
of predicted UARS telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
***** ALL PEM COMMUNICATION SHOULD NOW BE THROUGH THE PEM2K:: NODE *****
The PEM team reports that the PEM:: node computer has failed and has
been taken off line. PEM:: appears to have hardware problems that take the
system down every few hours. Automatic transfer NRT tests to PEM2K have
succeed however the NRT files have to be verified against those captured by
PEM to confirm the NRT process is correct.
PEM/AXIS was activated in its half power mode on June 24 and deactivated
on June 27. The PEM/AXIS was reactivated in full power mode on July 3. At
the current time, AXIS is not generating energy deposition. SwRI is looking
into this. Tests of AXIS production code on the rehost computer are continuing.
At the present time, PEM energy deposition from its spectrometers is not
being generated. Recent data configuration changes have affected PEM software
causing multiple failures. These failures are being reviewed by the PEM team.
In order to correct for new telemetry requirements, PEM will need to update
processing code with a patch delivery. The delivery patch may include changes
which would allow the PEM software to execute on a DEC Alpha machine. GSFC is
helping PEM with some DEC Alpha particulars which have caused PEM level 3
code to fail on DEC Alpha machines (software below level 3 executes on the
DEC Alpha).
Updated HEPS calibration parameters have been finalized and are ready for
delivery. Updated MEPS quality calibration files have been finalized and are
ready for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell,
have been fixed; his suggestions have been incorporated with in the PEM
production code. Both updated HEPS and MEPS calibration data, along with the
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
HRDI continues to function normally, making wind measurements during
the sunlit parts of orbits nearly every day, and measuring nighttime winds
when power is available.
***********
WINDII
***********
WINDII has commenced operations again. The first attempt, to
observe PMCs, was not successful, owing to low power conditions.
However, the instrument is fine, and the next observing period
is scheduled for July 13.
***********
MPG
***********
On 3 July 2001 the MPG participated in a meeting called
to discuss the status of software development efforts
underway so that UARS will be able to use three TDRS
(rather than two) spacecraft for communications.
Daily processing of instrument Daily Activity Plans
continues with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 23 June 2001 (Orbit 53476, GMT Day 174) through 29 June 2001
(Orbit 53580, GMT Day 180).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 174/2343 53490 Power OFF
PEM 177/2308 53535 PEM (Z/H) PWR OFF
ACRIM 177/2308 53535 POWER OFF
SUSIM 177/2308 53535 POWER OFF
PEM 178/2309 53550 AXIS-1 POWER OFF
SOLSTICE 179/2009 53563 Power OFF (LOW POWER)
A Reverse-to-Forward Yaw Around maneuver (number 100) was successfully executed
in Orbit 53558 beginning 179/10:54:41 and completing at 179/11:31:51 on 28 June
2001. The SOLSTICE instrument was in day only and all other instruments were
in their off or standby mode (including HRDI and PEM). Due to the SA position
at 269 degrees, the effective night length of the spacecraft after the maneuver
was approximately 57.07 minutes. Post maneuver analysis of data indicate that
during the Yaw Around, the end-of-night load bus voltage was 23.8 volts and the
average depths of discharge (DODs) on Batteries 2 and 3 were approximately 22.0
and 23.5 percent respectively. The next (forward-to-reverse) YAW is scheduled
for 8 August 2001 (DOY 220).
The beta angle ranged between the angles of 19.0 degrees to 1.7 degrees for
this report period. The beta angle is now increasing towards a maximum angle
of 78 degrees on 20 July 2001 (DOY 201).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.6 to 1.9 degrees C. Additional battery performance data for each
battery on 22 June 2001 and 29 June 2001 are provided in Appendix D.
The clock error ranged from Š14.5 to 12.8 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
„ OUT-OF-LIMITS = 00
„ ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
„ FOT = 09
„ HALOE = 01
„ SOLSTICE = 01
„ ACRIM = 01
„ SUSIM = 01
„ HRDI = 02
„ PEM = 02
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
-None this report period
One (1) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/316 - 2287.496493
· Transponder B: 01/119 - 2287.499116
GROUND SYSTEM OPERATION
Zero (0) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 01/179 show the observatory data loss to be 301 hours, 40
minutes, 44.840 seconds. This shows a loss of 38 minutes and 32.112 seconds
since the last report period. This is a 0.39 percent data loss, which equals a
99.61 percent data capture for the mission.
The increase of 38 minutes and 32.112 seconds occurred over two (2) days:
· 25.440 seconds on Day 175 due to a Data Hit.
· 38 minutes and 07.112 seconds on Day 179 due to Station preserving power
for YAW maneuver
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT continues to fine-tune the TAM alignment by loading new alignment
matrices. A new alignment is required for the reverse flight direction. This
is a slow process and it will take many alignments to lock in on the most
accurate solution.
An instrument anomaly was experienced on 05 May 2001. The instrument involved
was the ACRIM II instrument. All ACRIM telemetry for currents and voltages all
dropped to zero values suddenly. The ACRIM team in coordination with the
UARSFOT is conducting analysis and command tests to determine the cause of the
anomaly. Analysis shows that the anomaly was caused due to failed circuitry in
the ACRIM electronics box. The FOT is awaiting a future plan of action from
the ACRIM team.
MEETINGS
The FOT attended a UARS Monthly Status Meeting on Wednesday, June 27, 2001
The FOT attended a UARS SERB/TAC RS Meeting on Wednesday, June 27, 2001.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 08 August 2001 (DOY 220). This
maneuver will be a Forward to Reverse Flight maneuver (Southern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 29 June 2001
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
53479 174/0543 01-413 OBC Clk Adj -20.6 to 0.0 msec/day
53490 174/2343 01-414 HALOE Power OFF
53495 175/0808 01-416 OBC Clk Adj 0.0 to 20.6 msec/day
53510 176/0733 01-417 OBC/CLK Clk Adj 20.6 to 0.0 msec/day
53516 176/1649 01-418 OBC/CLK Clk Adj 0.0 to -20.6 msec/day
53535 177/2308 01-419 SOLSTICE TO DAY-ONLY OPS
53535 177/2308 01-420 PEM PEM (Z/H) PWR OFF
53535 177/2308 01-421 ACRIM POWER OFF
53535 177/2308 01-422 SUSIM POWER OFF
53535 177/2308 01-423 OBC/CLK Clk Adj -20.6 to 0.0 msec/day
53538 178/0405 01-424 NBTR NBTR PBK FOR FDF
53544 178/1428 01-425 HRDI THERMAL MICROLOADS
53550 178/2309 01-426 PEM AXIS-1 POWER OFF
53556 179/0925 01-427 OBC MONYAW & Table 37
53558 179/1135 01-427 MACS YAW CONTÕD
53558 179/1135 01-428 MACS TAM
53563 179/2009 01-429 SOLSTICE Power OFF (LOW POWER)
53572 180/1055 01-430 HRDI HRDI PB (INST) OFF
_____________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 29 June 2001
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 29 June 2001
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
53479 174/0506 Clk Adj -20.6 to 0.0 msec/day
53495 175/0817 Clk Adj 0.0 to 20.6 msec/day
53510 176/0736 Clk Adj 20.6 to 0.0 msec/day
53516 176/1652 Clk Adj 0.0 to -20.6 msec/day
53535 177/2322 Clk Adj -20.6 to 0.0 msec/day
___________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 29 June 2001
22 June 2001 - GMT Day 173 Beta = 21.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- -------- ---------- ------ ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/+0.0 12.0 0.000 0.0
2 +11.2/-112.0 +5.79/+4.51 +28.8/-14.4 24.5 1.040 20.8
3 +22.4/-56.0 +4.83/+3.23 +28.4/-14.4 24.3 1.040 20.5
29 June 2001 - GMT Day 180 Beta = 1.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- -------- ---------- ------ ------- ------
1 +728.0/+728.0 -3.05/-4.94 +0.0/+0.0 12.0 0.000 0.0
2 +11.2/-100.8 +5.47/+3.87 +28.0/-12.8 24.5 1.040 19.8
3 +22.4/-39.2 +4.51/+2.60 +27.2/-13.2 24.5 1.030 19.5
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
---- ---- ---------
-no changes V/T 5
_______________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 29 June 2001
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-033 FOT 53222 157/0220 N/A /CDEL function not working
01-034 FOT 53225 157/0754 N/A Unable to dump ATC
01-037 FOT 53292 161/1820 N/A Unable to load ATC
01-038 FOT 53454 172/1424 N/A TMON 1, 2 TRIPPED
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-032 FOT see list June 01 14170 Generic Late Acquisition
SEVNTFW01/26